OpenStudy (annah.bee):

Aye can someone help me with these problems?? So they're either A. infinite B. have no solution or C. One solution 1. 5x - 20 = 2(3x - 10) 2. 4x - 16 - 2x = 2(x - 8) 3. 6x - 15 = 3(2x - 6) 4. 5x - 3(x - 4 ) = 2 (x - 6) Thank you!!!!!

OpenStudy (precal):

5x-20=2(3x-10) 5x-20=6x-20

OpenStudy (annah.bee):

What I got was 1. one solution only 2. infinite 3. no solution 4. infinite

OpenStudy (precal):

5x-20=6x-20 -5x -5x -20=1x-20 +20 +20 0=x therefore one solution

OpenStudy (precal):

4x-16-2x=2(x-8) 2x-16=2(x-8) 2x-16=2x-16 notice both side are the same solution is infinitely many

OpenStudy (precal):

gotta go, hope that helps

OpenStudy (annah.bee):

Thank you so much!!!!!